OpenStudy (anonymous):

How do i solve the equation (6t+4)^2=3 ?

OpenStudy (lgbasallote):

square root both sides...tried that?

OpenStudy (anonymous):

(6t+4)^2=3 (6t+4)(6t+4)=3

OpenStudy (anonymous):

i squared but answer has a +- square root fraction

OpenStudy (lgbasallote):

square root both sides \[6t + 4 = \pm \sqrt 3\] subtract both sides by 4

OpenStudy (anonymous):

PEMDAS good ol' 6th grade. 1) Parentheses: they're as simplified as they can be 2) Exponent: you can do this (6t+4) * (6t+4) =3 3) Multiply: 36t^2+48t +16 = 3 got the rest from here?

OpenStudy (anonymous):

answer should be\[\left\{ \pm \sqrt{3-4}\right\}/6\]

OpenStudy (anonymous):

That's not correct. The correct answer is \[(-4\pm \sqrt{3})/6\]You can get this using the quadratic formula of \[(-b \pm \sqrt{b^{2}-4ac})/2a\]

OpenStudy (anonymous):

your answer is the one i see here but its the opposite way -4 is after square root 3 does it make a difference or is the same thing.

OpenStudy (anonymous):

As long as -4 is outside the root, it should be the same.

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

how do i end up with a square root of 3 and denominator of 6t?